# Artificial Intelligence Operating System; Low-Altitude Aviation; Autonomous Navigation; Multi-Sensor Fusion; Multi-Drone Coordination; Real-Time Task Management; Visual Processing; Ground Station Management

### v1

@@ Abstract
This paper introduces a comprehensive artificial intelligence operating system tailored for low-altitude aviation applications, integrating cutting-edge technologies for enhanced performance, safety, and efficiency. The system comprises six core components: OrinFlight OS, a high-performance operating system optimized for real-time task execution; UnitedVision, a versatile visual processing module supporting advanced image analysis; UnitedSense, a multi-sensor fusion module providing precise environmental modeling; UnitedNavigator, a dynamic path-planning and navigation system; UnitedMatrix, enabling multi-drone coordination and task execution; and UnitedInSight, a ground station for monitoring and management. Complemented by the UA DevKit low-code platform, the system facilitates user-friendly customization and application development. Leveraging NVIDIA Orin’s computational power and advanced AI algorithms, this system addresses complex challenges in modern aviation, offering robust solutions for navigation, perception, and collaborative operations. This work highlights the system’s architecture, features, and potential applications, demonstrating its ability to meet the demands of intelligent aviation environments.

### v2

@@ Abstract
This paper introduces a high-performance artificial intelligence operating system tailored for low-altitude aviation, designed to address key challenges such as real-time task execution, computational efficiency, and seamless modular collaboration. Built on a powerful hardware platform and leveraging the UNIX architecture, the system implements a distributed data processing strategy that ensures rapid and efficient synchronization across critical modules, including vision, navigation, and perception. By adopting dynamic resource management, it optimally allocates computational resources, such as CPU and GPU, based on task priority and workload, ensuring high performance for demanding tasks like real-time video processing and AI model inference. Furthermore, the system features an advanced interrupt handling mechanism that allows for quick responses to sudden environmental changes, such as obstacle detection, by prioritizing critical tasks, thus improving safety and mission success rates. Robust security measures, including data encryption, access control, and fault tolerance, ensure the system’s resilience against external threats and its ability to recover from potential hardware or software failures. Complementing these core features are modular components for image analysis, multi-sensor fusion, dynamic path planning, multi-drone coordination, and ground station monitoring. Additionally, a low-code development platform simplifies user customization, making the system adaptable to various mission-specific needs. This comprehensive approach ensures the system meets the evolving demands of intelligent aviation, providing a stable, efficient, and secure environment for complex drone operations.